OpenStudy (lgbasallote):

can someone refresh me on how to do this find the general solution using laplace transform y' = e^t ; y(0) = 2

OpenStudy (anonymous):

firstly apply laplace transform to both side of equation

OpenStudy (lgbasallote):

here's what i got.. \[y' = e^t\] \[\implies sy(s) - y'(0) = \frac{1}{s-a}\] \[\implies sy(s) - 2 = \frac{1}{s-a}\] \[\implies sy(s) = \frac{1 + 2(s-a)}{s-a}\] \[\implies y(s) = \frac{1 + 2s - 2a}{s(s-a)}\] idk if it's right...and even if it is idk how to proceed from there

OpenStudy (lgbasallote):

i think i did something wrong...

OpenStudy (anonymous):

? \[y(s)=\frac{1}{s(s-a)}+\frac{2}{s}\]

OpenStudy (anonymous):

\[L(y')=sy(s)-y(0)\]

OpenStudy (lgbasallote):

uhh yeah...lemme do this again

OpenStudy (anonymous):

and u have a=1

OpenStudy (anonymous):

you combined but its much easier to keep them seperate

OpenStudy (lgbasallote):

what do you mean i have a=1

OpenStudy (anonymous):

just use partial n the first and the second is more simple

OpenStudy (anonymous):

he means a=1 for \[e^{at}=e^{1t}\]

OpenStudy (lgbasallote):

ohh lol yeah...i forgot

OpenStudy (anonymous):

\[\frac{1}{s(s-1)}=\frac{A}{s}+\frac{B}{s-1}\] use coverup method

OpenStudy (lgbasallote):

coverup method?

OpenStudy (lgbasallote):

i am completely clueless about what you're doing now....although im getting the feeling you're getting this more complicated than it is...

OpenStudy (anonymous):

cover up method is just another way of partial fractions if you like to expand thats fine also \[1=(s-1)a+sB when s=0 -1=A when s=1 B=1

OpenStudy (lgbasallote):

did you tke the inverse laplace in the last step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i erased so that you could gt the answer yourself though

OpenStudy (anonymous):

you can check it at the end for the initial to make sure it's write

OpenStudy (anonymous):

right*

OpenStudy (lgbasallote):

ohhh so you have to take tehe inverse laaplace in the end to get the final answer? always?

OpenStudy (anonymous):

yes because you have \[F(s)\] and you want \[f(t)\]

OpenStudy (lgbasallote):

ohhh

OpenStudy (lgbasallote):

i didnt know tha..thanks

OpenStudy (anonymous):

\[L({f(t)})=F(s)\] \[L^{-1}(F(s))=f(t)\]

OpenStudy (anonymous):

have you taken linear... these are a lot like kernels the chapter about them... my book actually classifies them as kernel (K(s))

OpenStudy (lgbasallote):

i dont believe so